What s the minimom ini hal veloaty a rockst must be firea t ablk 000 km oway . VsC angle for mar raye
Answer #1

Range in projectile motion is given by:

R = (V0^2*sin 2\theta)/g

Now for minimum velocity, projectile angle should be 45 deg, So that range can be maximum

\theta = 45 deg

sin 2\theta = sin (2*45 deg) = sin 90 deg = 1

V0 = initial velocity of projectile

g = gravitational acceleration = 9.81 m/sec^2

R = range of rocket = 1000 km = 10^6 m

So,

V0^2 = g*R/sin 2\theta = g*R

V0 = sqrt (g*R)

V0 = sqrt (9.81*10^6)

V0 = 3132.1 m/sec = minimum velocity of rocket

Please Upvote.